OpenStudy (piercetheveil47):

What is the exact distance from (−1, 4) to (6, −2)? square root of 80. units square root of 82. units square root of 85. units square root of 89. units

OpenStudy (piercetheveil47):

is it B?

OpenStudy (triciaal):

Do you know the pythagorean theorem? this is the same as the distance formula. the distance between 2 points is the sqrt of

OpenStudy (triciaal):

sum of (the change in y )^2 and the (change in x)^2

OpenStudy (piercetheveil47):

so its not B?

OpenStudy (piercetheveil47):

jk its C :)

OpenStudy (triciaal):

your 2 points are (−1, 4) and (6, −2) the change in y = -2 -4 = -6 the change in x = 6--1= 7 distance = sq rt of (36 + 49) = sq rt of 85

OpenStudy (piercetheveil47):

thanks!

OpenStudy (triciaal):

yes C you are welcome